OpenStudy (anonymous):

does anybody know what questions the flvs teachers ask for algebra 1 modules 1-5???

OpenStudy (anonymous):

for grade 9??

Mehek (mehek14):

they ask about the whole Module to see what you learned I'm pretty sure each teacher asks different questions One of the question I was asked was to make an equation and solve it

Mehek (mehek14):

She also asked the meaning of the terms and how to use PEMDAS

Mehek (mehek14):

Also this: add- plus, more than, increase by, sum of, total, sub-less than, decreased by, multiply- product, doubles, times, triples, divide- 1/3, 1/2 ratio, quotient. 26/30

OpenStudy (anonymous):

i need equations then because i wouldnt know what to give them. im not good with phone calls at all. i get really nervous and forget everything i have learned and its super scary and frustrating what do i say

Mehek (mehek14):

I'm just saying your teacher COULD ask you every teacher probably asks different questions she could give you an equation to solve instead

OpenStudy (anonymous):

okay thanks but can you give me an equation real quick and help me solve it. im having trouble anyway

Mehek (mehek14):

Simple: 2x - 10 = 6

Mehek (mehek14):

2x - 10 = 6 first add 10 on both sides 2x = 16 divide 2 on both sides x = 8

OpenStudy (anonymous):

what about the ones that look like these??3x - 2y less than or greater to 4

Mehek (mehek14):

that's in module 1?

OpenStudy (anonymous):

5x + 2y ≥ 13

OpenStudy (anonymous):

those are in module 5

OpenStudy (anonymous):

can you help show me how to solve them

Mehek (mehek14):

Well right now practice for 1

Mehek (mehek14):

they should explain how to solve it in the module

OpenStudy (anonymous):

Delilah I have been a connections academy student for six years, I haven't been in FLVS for 3 years, but have you webmailed your teacher? They typically will provide you with the modules.

OpenStudy (anonymous):

first add 10 on both sides 2x = 16 divide 2 on both sides x = 8 what were the both sides

Mehek (mehek14):

2x - 10 = 6 +10 +10 both sides of the equal sign 2x = 6 -- --- 2 2 x = 8

OpenStudy (anonymous):

what if it was 3x+ 10 = 6??

Mehek (mehek14):

3x + 10 = 6 subtract 10 3x = -4 divide by 3 x = -1.33

OpenStudy (anonymous):

oh my gosh thank you. youre helping

OpenStudy (anonymous):

can you help me now with these |x + 2| + 16 = 14

OpenStudy (anonymous):

i need help :( |2x + 6| − 4 = 20

OpenStudy (anonymous):

@mehek14

OpenStudy (anonymous):

HELLOOO?????

Mehek (mehek14):

positive sign 2x + 6 - 4 = 20 2x +2 = 20 2x = 18 x = 9 negativesign -2x - 6 - 4 = 20 -2x -10 = 20 -2x =30 x = -15